OpenStudy (anonymous):

solve the system by substitution -4.5x-2y=-12.5 3.25x=-0.75 please help, I have to show it step by step and I have no idea what i'm doing :/

OpenStudy (anonymous):

Sorry messed up its -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

Solve for one of the variables. In this case solving for y in the second equation would be quickest.

OpenStudy (anonymous):

y=0.75+3.25x

OpenStudy (anonymous):

substitute (0.75+3.25x) for y in the first equation. -4.5x-2(0.75+3.25x)=-12.5 solve for x and plug that value into either of the original equations to find y

OpenStudy (anonymous):

Thank you :)